OpenStudy (erinweeks):

Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0

OpenStudy (anonymous):

yo u go to james madison? which one? my cousin goes there too

OpenStudy (erinweeks):

A. \[i \sqrt{11}, -i \sqrt{11}, -5\] B. \[\sqrt{11}, -\sqrt{11}, 5\] C. \[i \sqrt{11}, -5\] D. \[i \sqrt{11}, -i \sqrt{11}, \]

OpenStudy (campbell_st):

start with x = 5 to test it find f(5) if its equal to zero then (x - 5) is a factor

OpenStudy (erinweeks):

Yes i go to James Madison! I do online

OpenStudy (erinweeks):

how do i find f(5)?

OpenStudy (campbell_st):

then if (x -5) is a factor then use polynomial division to to find the quadratic factor... then solve the quadratic

OpenStudy (campbell_st):

substitute 5 into the equation in place of x.... this is the factor theorem if f(5) = 0 ... then (x - 5) is a factor

OpenStudy (campbell_st):

but given the answer choices ... there is only 1 with a root x = 5 which makes it easy

OpenStudy (erinweeks):

i got B?

OpenStudy (campbell_st):

makes sense to me... but its interesting that you don't know about the factor theorem... and you need to look up the rational roots theorem...

jimthompson5910 (jim_thompson5910):

alternative way -x^3 + 5x^2 - 11x + 55 = 0 (-x^3 + 5x^2) + ( -11x + 55) = 0 -x^2(x - 5) + ( -11x + 55) = 0 -x^2(x - 5) - 11(x - 5) = 0 (-x^2 - 11)(x - 5) = 0 -x^2 - 11 = 0 or x - 5 = 0 x^2 + 11 = 0 or x - 5 = 0 x^2 = -11 or x = 5 x = sqrt(-11), x = -sqrt(-11), or x = 5 x = i*sqrt(11), x = -i*sqrt(11), or x = 5

OpenStudy (anonymous):

The possible roots of a polynomial are the factors of C (constant) divided by the factors of A (coefficient of the first term). So the possible roots are: +1,-1,+5,-5,+55,-55,+11,-11 For any of these to be the roots, the function must equal to zero when we substitute the value for x. We try +5 and it gives us a zero; f(5) = -(5)^3 + 5(5)^2 - 11(5) + 55 = 0 Therefore, x-5 is a factor. Now we can do long division and divide the original function by x-5. (-x^3 + 5x^2 - 11x + 55) / x-5 = -x^2 - 11 (Through long division) Now we can re-write our original function as f(x) = (x - 5)(-x^2 - 11) Now we factor the quadratic part to obtain additional roots. Note these roots will be imaginary as this quadratic has no real roots. -x^2 - 11 = 0 -x^2 = 11 x = sqrt(-11) or -sqrt(-11) We can rewrite each root in terms f "i" like this: x = i*sqrt(11), -i*sqrt(11), 5